OpenStudy (anonymous):

What is the prime factorization of 2025

OpenStudy (anonymous):

The Prime Factors are: 3 x 3 x 3 x 3 x 5 x 5 http://www.calculatorsoup .com/calculators/math/prime-factors.php use this link

jigglypuff314 (jigglypuff314):

Hello @ellarocks1 and Welcome to OpenStudy! :) to get the prime factorization of a number, keep dividing by numbers you know can be divided into 2025 for example if you want the prime factorization of 36 36 / 2 = 18 18 / 2 = 9 9 / 3 = 3 <- which is prime so you have two 2s and two 3s so the prime factorization of 36 would be 2^2 * 3^2